OpenStudy (anonymous):

for exponential function of the form f(x)=a^x which of the following expressions is equal to f(r).f(s) ? (A) f(r.s) (B) f (r^s) (C) f(r+s)

OpenStudy (anonymous):

\[f(r).f(s)=a^r . a^s=a^{r+s}\]

OpenStudy (anonymous):

i dont understand !!! so what would the answer be??

OpenStudy (anonymous):

note that\[f(r+s)=a^{r+s}\]

OpenStudy (anonymous):

soo it would be f(r^s)

OpenStudy (anonymous):

omg thx u i got it rite!!! ;)